# Entanglement Entropy for Screened Interactions via Dimensional Mapping to Harmonic Oscillators

### v1

@@ V Perturbative evaluation of entanglement entropy
which induced the corresponding expansion of the reduced density eigenvalue,
@@
Zeroth-order entropy. At leading order, the reduced density eigenvalue is purely Gaussian,
@@
Gaussian contribution. The Gaussian part originates entirely from the linear shift
@@
Non-Gaussian contribution. The remaining part, $\rho^{(1)}_{\mathrm{NG}}(k)$, arises from genuine interaction effects. However, at first order the reduced density matrix is normalized such that
@@
Gaussian contribution. The Gaussian part originates solely from the second-order expansion of the width parameter $\beta(\alpha)$ and corresponds to a further renormalization of the harmonic oscillator scale. Evaluating the integrals in Eq.(89) using $\rho^{(2)}_{\mathrm{G}}(k)$ yields
@@
Non-Gaussian contribution. The genuinely interaction-induced correction arises from the $\rho^{4}$ perturbation, which produces polynomial deformations of the Gaussian kernel of the reduced density matrix. At this order, the non-Gaussian part takes the schematic form
@@
VI Structure and Rules of the $\rho^{2n}$ Perturbations
@@ A.1 Separated wavefunction
$\frac{\partial\Psi}{\partial x_{1}}=\frac{\partial\Psi}{\partial x_{c}}\frac{\partial x_{c}}{\partial x_{1}}+\frac{\partial\Psi}{\partial x}\frac{\partial x}{\partial x_{1}}=\frac{m_{1}}{M}\frac{\partial\Psi}{\partial x_{c}}+\frac{\partial\Psi}{\partial x}$ (103)
@@
$\frac{\partial\Psi}{\partial x_{2}}=\frac{\partial\Psi}{\partial x_{c}}\frac{\partial x_{c}}{\partial x_{2}}+\frac{\partial\Psi}{\partial x}\frac{\partial x}{\partial x_{2}}=\frac{m_{2}}{M}\frac{\partial\Psi}{\partial x_{c}}-\frac{\partial\Psi}{\partial x}$ (104)
@@
$\frac{\partial^{2}\Psi}{\partial x_{1}^{2}}=\frac{m_{1}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}+\frac{\partial^{2}\Psi}{\partial x^{2}}$ (105)
@@
$\frac{\partial^{2}\Psi}{\partial x_{2}^{2}}=\frac{m_{2}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}+\frac{\partial^{2}\Psi}{\partial x^{2}}$ (106)
@@
$-\frac{\hbar^{2}}{2m_{1}}\frac{m_{1}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}-\frac{\hbar^{2}}{2m_{1}}\frac{\partial^{2}\Psi}{\partial x^{2}}-\frac{\hbar^{2}}{2m_{2}}\frac{m_{2}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}$
$-\frac{\hbar^{2}}{2m_{1}}\frac{\partial^{2}\Psi}{\partial x^{2}}-g^{2}\frac{e^{-\alpha m|x|}}{|x|}\Psi=-B\Psi$ (107)
@@ A.2 Explicit calculation for $f(\rho)$
$\frac{df}{dx}=\frac{d\rho}{dx}\frac{df}{d\rho}=\frac{1}{2\sqrt{x}}\frac{df}{d\rho}=\frac{1}{2\rho}\frac{df}{d\rho}$ (109)
@@
$\frac{d^{2}f}{dx^{2}}=\frac{d}{dx}\left(\frac{1}{2\sqrt{x}}\frac{df}{d\rho}\right)=-\frac{1}{4x^{3/2}}\frac{df}{d\rho}+\frac{1}{2\sqrt{x}}\frac{df}{dx}\frac{d^{2}f}{d\rho^{2}}$ (110)
@@
$\frac{d^{2}f}{dx^{2}}=-\frac{1}{4\rho^{3}}\frac{df}{d\rho}+\frac{1}{4\rho^{2}}\frac{d^{2}f}{d\rho^{2}}$ (111)
@@
$\frac{d^{2}f}{d\rho^{2}}-\frac{1}{\rho}\frac{df}{d\rho}=Ah(\rho)\frac{d^{2}\phi}{d\rho^{2}}-C\frac{h(\rho)\phi(\rho)}{\rho^{2}}$ (113)
$4B_{r}\rho^{2}f=\frac{1}{2}\mu\,\omega^{2}\rho^{2}h(\rho)\phi(\rho)\\$ (114)
$4g^{2}e^{-\alpha m\rho^{2}}f=Eh(\rho)\phi(\rho)$ (115)
@@
Here, the second condition Eq.(114) demands that $f=h(\rho)\phi(\rho)$ whereas the third condition Eq.(115) demands that $f=e^{\alpha m\rho^{2}}h(\rho)\phi(\rho)$. We can try and redefine $f$ such that $f=j(\rho)\phi(\rho)$, $j(\rho)=e^{\alpha m\rho^{2}}h(\rho)$
We intend to find the form of $h(\rho)$ so that we can use it to transform Eq.(12) into Eq.(II). Second order derivative of $f$ is,
$\frac{df}{d\rho}=2\alpha m\rho e^{\alpha m\rho^{2}}h\phi+e^{\alpha m\rho^{2}}h^{\prime}\phi+e^{\alpha m\rho^{2}}h\phi^{\prime}$ (116)
$\frac{d^{2}f}{d\rho^{2}}$ $=2\alpha me^{\alpha m\rho^{2}}h\phi+4\alpha^{2}m^{2}\rho^{2}e^{\alpha m\rho^{2}}h\phi$
$+2\alpha m\rho e^{\alpha m\rho^{2}}h^{\prime}\phi+2\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}$
$+2\alpha m\rho e^{\alpha m\rho^{2}}h^{\prime}\phi+e^{\alpha m\rho^{2}}h^{\prime\prime}\phi$
$+e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}+2\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}$
$+e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}+e^{\alpha m\rho^{2}}h\phi^{\prime\prime}$ (117)
Substituting Eq.(116) and (A.2) in L.H.S of condition 1, Eq.(113), we obtain
$4\alpha^{2}m^{2}\rho^{2}e^{\alpha m\rho^{2}}h\phi+4\alpha m\rho e^{\alpha m\rho^{2}}h^{\prime}\phi$
$+4\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}+2e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}$
$+e^{\alpha m\rho^{2}}h^{\prime\prime}\phi+e^{\alpha m\rho^{2}}h\phi^{\prime\prime}$
$-\frac{e^{\alpha m\rho^{2}}h^{\prime}\phi}{\rho}-\frac{e^{\alpha m\rho^{2}}h\phi^{\prime}}{\rho}$ (118)
@@
$4\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}+2e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}-\frac{e^{\alpha m\rho^{2}}h\phi^{\prime}}{\rho}=0$ (119)
@@
$4\alpha m\rho h+2\frac{dh}{d\rho}-\frac{h}{\rho}=0$ (120)
@@
$\int\frac{dh}{h}=\frac{1}{2}\int\frac{d\rho}{\rho}-\int 2\alpha m\rho d\rho$ (121)
@@
$f=A\sqrt{\rho}\phi(\rho)$ (123)
This form of $f(\rho)$ satisfies the relative coordinates Schrodinger equation for harmonic oscillator, yet, in our case, Eqs.(113)-(114) demand the use of pertubative expansion to obtain eigenvalue equation to correctly.

### v2

@@ 5 Perturbative evaluation of entanglement entropy
which induced the corresponding expansion of the reduced operator eigenvalue,
@@
Zeroth-order entropy. At leading order, the reduced density operator eigenvalue is purely Gaussian,
@@
Gaussian contribution: The Gaussian part originates entirely from the linear shift
@@
Non-Gaussian contribution: The remaining part, $\rho^{(1)}_{\mathrm{NG}}(k)$, arises from genuine interaction effects. However, at first order the reduced density matrix is normalized such that
@@
Gaussian contribution: The Gaussian part originates solely from the second-order expansion of the width parameter $\beta(\alpha)$ and corresponds to a further renormalization of the harmonic oscillator scale. Evaluating the integrals in eq.(78) using $\rho^{(2)}_{\mathrm{G}}(k)$ yields
@@
Non-Gaussian contribution: The genuinely interaction-induced correction arises from the $\rho^{4}$ perturbation, which produces polynomial deformations of the Gaussian kernel of the reduced density matrix. At this order, the non-Gaussian part takes the schematic form
@@
6 Structure and Rules of the $\rho^{2n}$ Perturbations
@@ A.1 Separated wavefunction
$\frac{\partial\Psi}{\partial x_{1}}=\frac{\partial\Psi}{\partial x_{c}}\frac{\partial x_{c}}{\partial x_{1}}+\frac{\partial\Psi}{\partial x}\frac{\partial x}{\partial x_{1}}=\frac{m_{1}}{M}\frac{\partial\Psi}{\partial x_{c}}+\frac{\partial\Psi}{\partial x}$ (A.1)
@@
$\frac{\partial\Psi}{\partial x_{2}}=\frac{\partial\Psi}{\partial x_{c}}\frac{\partial x_{c}}{\partial x_{2}}+\frac{\partial\Psi}{\partial x}\frac{\partial x}{\partial x_{2}}=\frac{m_{2}}{M}\frac{\partial\Psi}{\partial x_{c}}-\frac{\partial\Psi}{\partial x}$ (A.2)
@@
$\frac{\partial^{2}\Psi}{\partial x_{1}^{2}}=\frac{m_{1}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}+\frac{\partial^{2}\Psi}{\partial x^{2}}$ (A.3)
@@
$\frac{\partial^{2}\Psi}{\partial x_{2}^{2}}=\frac{m_{2}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}+\frac{\partial^{2}\Psi}{\partial x^{2}}$ (A.4)
@@
$-\frac{\hbar^{2}}{2m_{1}}\frac{m_{1}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}-\frac{\hbar^{2}}{2m_{1}}\frac{\partial^{2}\Psi}{\partial x^{2}}-\frac{\hbar^{2}}{2m_{2}}\frac{m_{2}^{2}}{M^{2}}\frac{\partial^{2}\Psi}{\partial x_{c}^{2}}-\frac{\hbar^{2}}{2m_{1}}\frac{\partial^{2}\Psi}{\partial x^{2}}-g^{2}\frac{e^{-\alpha m|x|}}{|x|}\Psi=-B\Psi$ (A.5)
@@ A.2 Explicit calculation for $f(\rho)$
$\frac{df}{dx}=\frac{d\rho}{dx}\frac{df}{d\rho}=\frac{1}{2\sqrt{x}}\frac{df}{d\rho}=\frac{1}{2\rho}\frac{df}{d\rho}$ (A.7)
@@
$\frac{d^{2}f}{dx^{2}}=\frac{d}{dx}\left(\frac{1}{2\sqrt{x}}\frac{df}{d\rho}\right)=-\frac{1}{4x^{3/2}}\frac{df}{d\rho}+\frac{1}{2\sqrt{x}}\frac{df}{dx}\frac{d^{2}f}{d\rho^{2}}$
@@
$\frac{d^{2}f}{dx^{2}}=-\frac{1}{4\rho^{3}}\frac{df}{d\rho}+\frac{1}{4\rho^{2}}\frac{d^{2}f}{d\rho^{2}}$ (A.8)
@@
$\frac{d^{2}f}{d\rho^{2}}-\frac{1}{\rho}\frac{df}{d\rho}=Ah(\rho)\frac{d^{2}\phi}{d\rho^{2}}-C\frac{h(\rho)\phi(\rho)}{\rho^{2}}$ (A.10)
$4B_{r}\rho^{2}f=\frac{1}{2}\mu\,\omega^{2}\rho^{2}h(\rho)\phi(\rho)\\$ (A.11)
$4g^{2}e^{-\alpha m\rho^{2}}f=Eh(\rho)\phi(\rho)$ (A.12)
@@
Here, the second condition eq.(A.11) demands that $f=h(\rho)\phi(\rho)$ whereas the third condition eq.(A.12) demands that $f=e^{\alpha m\rho^{2}}h(\rho)\phi(\rho)$. We can try and redefine $f$ such that $f=j(\rho)\phi(\rho)$, $j(\rho)=e^{\alpha m\rho^{2}}h(\rho)$. We intend to find the form of $h(\rho)$ so that we can use it to transform eq.(11) into eq.(12). Second order derivative of $f$ is,
$\frac{df}{d\rho}=2\alpha m\rho e^{\alpha m\rho^{2}}h\phi+e^{\alpha m\rho^{2}}h^{\prime}\phi+e^{\alpha m\rho^{2}}h\phi^{\prime}$ (A.13)
$\frac{d^{2}f}{d\rho^{2}}$ $=2\alpha me^{\alpha m\rho^{2}}h\phi+4\alpha^{2}m^{2}\rho^{2}e^{\alpha m\rho^{2}}h\phi+2\alpha m\rho e^{\alpha m\rho^{2}}h^{\prime}\phi$
$+2\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}+2\alpha m\rho e^{\alpha m\rho^{2}}h^{\prime}\phi+e^{\alpha m\rho^{2}}h^{\prime\prime}\phi$
$+e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}+2\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}+e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}+e^{\alpha m\rho^{2}}h\phi^{\prime\prime}$ (A.14)
Substituting eqs.(A.13) and (A.2) in l.h.s of condition 1 in eq.(A.10), we obtain
$4\alpha^{2}m^{2}\rho^{2}e^{\alpha m\rho^{2}}h\phi+4\alpha m\rho e^{\alpha m\rho^{2}}h^{\prime}\phi+4\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}$
$+2e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}+e^{\alpha m\rho^{2}}h^{\prime\prime}\phi+e^{\alpha m\rho^{2}}h\phi^{\prime\prime}-\frac{e^{\alpha m\rho^{2}}h^{\prime}\phi}{\rho}-\frac{e^{\alpha m\rho^{2}}h\phi^{\prime}}{\rho}$ (A.15)
@@
$4\alpha m\rho e^{\alpha m\rho^{2}}h\phi^{\prime}+2e^{\alpha m\rho^{2}}h^{\prime}\phi^{\prime}-\frac{e^{\alpha m\rho^{2}}h\phi^{\prime}}{\rho}=0$ (A.16)
@@
$4\alpha m\rho h+2\frac{dh}{d\rho}-\frac{h}{\rho}=0$ (A.17)
@@
$\int\frac{dh}{h}=\frac{1}{2}\int\frac{d\rho}{\rho}-\int 2\alpha m\rho d\rho$ (A.18)
@@
$f=A\sqrt{\rho}\phi(\rho)$ (A.20)
This form of $f(\rho)$ satisfies the relative coordinates Schrodinger equation for harmonic oscillator, yet, in our case, eqs.(A.10)-(A.11) demand the use of pertubative expansion to obtain eigenvalue equation.